# Estimating Parameter Fields in Multi-Physics PDEs from Scarce Measurements

### v1

@@ II Thermal Runaway in Batteries
Measurements: Considering real-life constraints, only surface temperatures can be measured, and only at limited locations. For training, we take just three measurement locations, with 15 spatial-temporal measurements each, during the initial 100 minutes before the irreversible temperature rise at around 150 minutes in TR.
Analysis: Despite this limitation, Neptune accurately estimates the field parameters ($C_{p},k_{x}$, and $k_{z}$). For the prismatic battery, the estimated parameters (Fig. 3b) show MAE of 1.88, 0.052, and 0.001, compared to reference values of 1400, 13, and 0.8, respectively. For the cylindrical battery, the scalar parameter $A_{e}$ is jointly estimated, with an estimated value of $5.19\times 10^{25}$ showing a low percentage error of 0.97% compared to the reference $A_{e}=5.4\times 10^{25}$. Similar accuracy is achieved for other parameters, with results detailed in Supplementary Note 2. Furthermore, Neptune also demonstrates robustness under noisy measurements; for the prismatic battery, results for noise levels of 1% and 10% are detailed in Table 1.
@@
Figure 4: Field parameter estimation for flow in porous media. a. A demonstration of flow phenomena, and random measurements are taken as the ground truth for model training. b. The estimated hydraulic conductivity $K$ compared to the reference $K$. c. The estimated field parameter $K$ and its estimation error compared to the reference, for different numbers of training samples. The magnitudes are normalized. d. The particle concentration predictions (state responses) at different times. The predictions align well with the ground truth, even for extrapolation at $t=20$ minutes. e. The parameter estimation performance compared to different numbers of training samples. Error bands are derived from standard deviations from multiple estimations and represent the variability in the performance. f. The response prediction performance compared to different numbers of training samples.
@@ III Flow in Porous Media
Problem Setup: A 2D computational domain of $L_{x}=1m$ and $L_{y}=0.5m$ is defined [92], as shown in Fig.4a. Using just 32 random spatial-temporal concentration measurements $u$ over 0–16 minutes, the field parameter $K$ can be robustly estimated. Since the right-hand side of equation 7 is 0, multiple solutions exist, so the estimated $K$ is normalized and compared to the normalized reference (see Fig. 4b).
@@
To demonstrate the robustness of Neptune, parameter estimation is repeated with varying sample sizes. Figure 4e shows the parameter estimation performance, while Fig. 4f displays the corresponding response prediction accuracy, with errors significantly reduced at 64 samples. Ablation studies on neural network sizes, detailed in Supplementary Note 4, further validate the model’s robustness and generalizability. Additionally, a spline-based parameter fitting method was tested as a non-neural network strategy but showed lower performance than Neptune (around 2 times larger mean errors in both the parameter estimation and response prediction). The spline-based parameter fitting setups and results can be found in Supplementary Note 6.
@@
Figure 5: Parameter estimation in cell migration and proliferation. a. Parameter estimation results and comparisons. Neptune assumes all three parameters are density-dependent. The results show an increasing $\gamma$ value as density increases, while the other two parameter estimations exhibit minimal variation and can be considered as constants. This aligns well with assumptions from previous research [95, 96]. b. Density response predictions. Using the estimated field parameters, the density response at four different timesteps is displayed. Compared to results using the reference method (PINN-SR) and the scalar parameters setting, Neptune more accurately matches the experimental data at 36-hour and 48-hour, particularly in the regions with higher density values on both sides.
@@ IV Cell Migration and Proliferation
Analysis: As shown in the first graph of Fig. 5a, Neptune estimates an increasing $\gamma$ with cell density, consistent with previous research [97, 98]. Field estimations also suggest $\lambda_{1}$ and $\lambda_{2}$ are likely scalars. We also assume space-dependent parameters for estimation: results showing parameter distributions with high (peak) values at two boundaries, demonstrating that the field parameters are density (or state) dependent. In the third graph, while the field estimation for $\lambda_{2}$ shows large values at low densities and stabilizes as density increases, this behavior is primarily due to highly noisy experimental data in regions with low cell numbers, rather than being attributed to the nature of the field parameter itself. In contrast, PINN-SR, limited by its methodology, cannot estimate field parameters, and its scalar estimations show slightly different combinations but remain within a reasonable range. This limitation is evident at $12h$ (first graph) and $48h$ (last graph), where it fails to capture enhanced cell spreading.
Accurate field parameter estimation excels in predicting cell density distributions over time, as shown in Fig. 5b. For such sparse and noisy data, PINN-SR achieves a median error of $6.53\times 10^{-5}$. In contrast, Neptune, assuming density-dependent parameters, achieves the highest accuracy with an error of $4.71\times 10^{-5}$. Across experimental data with varying initial cell densities [13], Neptune achieves 28-47% lower median errors. Additional results are detailed in Supplementary Note 2.
@@ V Discussion
In this paper, we presented Neptune for accurate field parameter estimation under scarce observation data, demonstrated by various PDE problems in emerging fields where physical parameters are unknown and important. Specifically, we estimated the thermal conductivity and heat capacity in coupled thermal problems, enabling accurate TR prediction. In porous media flow, spatially varying hydraulic conductivity was recovered, improving the modeling of phenomena such as contaminant transport. For cardiac EP, tissue electrical conductivity was inferred solely from voltage measurements, offering potential for detecting fibrosis and other localized pathologies. Finally, the cell migration and proliferation case demonstrates effective parameter estimation under real-world noisy and sparse data conditions. It also highlights the method’s capability in discovering governing equations and its generalizability in estimating scalar variables.
@@
In terms of utility, Neptune demonstrates its generalizability and robustness to estimate multiple parameters simultaneously across various dependencies (spatial, state, and even time), each exhibits distinct patterns. The proposed two-stage learning strategy further enhances effectiveness by handling parameters across different scales. By first estimating scalar magnitudes, it significantly reduces the search space, allowing neural networks to focus on learning local variations within the parameter field. Furthermore, we extend this discussion to demonstrate the superior performance of our strategy compared to other training approaches, including: (1) estimating only scalars, (2) directly estimating each field parameter in one step using separate neural networks; (3) estimating all field parameters within a single neural network; and (4) jointly estimating scalar magnitudes and field variations in a single step. Detailed comparisons are provided in Supplementary Note 5. Compared to other fitting methods, Neptune significantly outperforms spline-based approaches in capturing complex parameter distributions. In addition, the use of FDM makes the method easily adaptable to a wide range of PDEs. While less effective for handling shock waves or large discontinuities, it offers favorable computational efficiency and supports gradient-based optimization.
@@
Compared to sparse identification methods like PINN-SR, Neptune models full field parameter distributions rather than scalars, achieving 27–47% improvement in forward response accuracy. This methodological shift enables more detailed modeling of the underlying physical processes, leading to improved predictive accuracy. Additionally, compared to PINN, our framework offers stronger temporal extrapolation, better adherence to physical laws, and greater adaptability to boundary and initial condition changes. It also requires less data, making it ideal when measurements are limited or costly. However, inverse problems like parameter estimation remain ill-posed and prone to local minima, especially with complex fields or limited data. These challenges can be mitigated by using informed initial guesses and incorporating prior knowledge, such as parameter magnitudes, smoothness penalties, and strategic data selection. Future work may extend the method to more general finite element frameworks and explore its application to modeling and parameter estimation in chaotic systems.
@@ VI Methods
Field parameter modeling with neural networks Neptune models a unknown field parameter $p$ as:
@@
where $\tilde{p}$ denotes a learnable scalar and $\mathcal{N}$ represents the neural network with $m$ layers and learnable weights as $\bm{\theta}$. A detailed architecture with skip connections can be referred to Fig. 6. $f$ denotes a neural layer and $\circ$ represents function composition. The input to $\mathcal{N}$ can be discretized values of space $\mathcal{X}$ or state $\mathcal{S}$, and are properly scaled (i.e., -1 to 1 from $\mathcal{X}$ or observed $\mathcal{S}$). Leaky ReLU activation functions are used in all layers except the final one, which omits activation functions. For multiple field parameters $\tilde{\mathcal{P}}=[p_{1},p_{2},\dots]$, Neptune simply employs different scalars $\tilde{p}_{i}$ and networks $\mathcal{N}_{{\bm{\theta}}_{i}}$ for parallel modeling, where $i=1,2,\dots$
Figure 6: The schematic diagram of the two-stage field parameter estimation strategy and the neural network architecture. a. The feed-forward neural network. Left: overall architecture. Right: residual block structure. For flow and cardiac electrophysiology problems, the input to the network is the spatial coordinate $\mathcal{S}$. For battery thermal runaway and cell migration, the input is the state variable $\mathcal{X}$ (e.g., $T$ or $\rho$), with fewer layers used. The network output $\mathcal{N}_{{\bm{\theta}}}$ is scaled to the physical parameter as $\tilde{\mathcal{P}}\big{(}1+\mathcal{N}_{{\bm{\theta}}}\big{)}$. The reference value $\tilde{\mathcal{P}}$ is application-specific and determined via a two-step training strategy. b, Parameter estimation diagram. The left part illustrates the two-stage estimation for scalar and field parameters, while the right part shows the numerical solving and optimization procedure.
@@
During the optimization, scalars $\tilde{\mathcal{P}}$ act as a physics-based prior and implicit regularizer, ensuring early physical stability of training and providing a well-scaled baseline for convenient expansion to field parameters in the second stage. Even if $\tilde{\mathcal{P}}$ converges with high error, its combination with the PDE remains physically meaningful and stable during optimization.
@@
Compared to directly learning $\tilde{\mathcal{P}}$ with a neural network, the first-stage scalar learning significantly accelerates training, as numerically solving the PDE with field parameter modeling is several times more expensive, depending on model complexity. It also reduces the search space and greatly stabilizes the second-stage learning of $\mathcal{N}_{\bm{\theta}}$.
@@
Neural network architecture and problem-specific design The feed-forward neural networks in the second stage are adaptively designed, with the number of residual blocks and neurons per layer customized for each application. Specifically, in the thermal battery problem, the scalar references for thermal conductivity in different directions ($\tilde{k}_{x}$, $\tilde{k}_{z}$) and heat capacity ($\tilde{C_{p}}$) are first determined. In the cell migration problem, three scalar parameters, $\tilde{\gamma}$, $\tilde{\lambda_{1}}$, and $\tilde{\lambda_{2}}$, are similarly estimated. In both cases, the neural network models state-dependent field parameters based on a single state variable, with a network input size of 1 (e.g., temperature $T$ or density $\rho$, rescaled approximately from -1 to 1). A compact network with 20 neurons per hidden layer and two residual blocks is used to ensure efficient training.
@@
Additionally, Supplementary Note 4 presents an ablation study evaluating the impact of neural network size on parameter estimation performance, while Supplementary Note 5 provides a detailed comparison of this training strategy against alternatives, including direct field parameter learning and using a single neural network for all parameters.
@@
During training, observations are split into mini-batches of size 16 (reduced to 8 for the cell migration problem due to limited samples). In the first stage, the loss is minimized to directly learn $\tilde{\mathcal{P}}$, with the learning rate adaptively decreasing from 1 to $10^{-2}$ until convergence, where $\tilde{\mathcal{P}}$ stabilizes without further change. Dashed lines in the figure illustrate the gradient flow during optimization. In the second stage, the loss is minimized to optimize the neural network weights $\bm{\theta}$, with a smaller learning rate ranging from $10^{-4}$ to $10^{-2}$ depending on network size. The best-performing $\bm{\theta}$, corresponding to the lowest loss value, is saved as the final trained model. The field parameter is then predicted using equation 14 with the corresponding spatial or state inputs.
@@
Importantly, Neptune differs fundamentally from traditional neural ODE approaches: instead of modeling unknown or partially known governing expressions with neural networks, it tackles complex PDEs with fully known equations, while neural networks serve only as correction terms for modeling unknown field parameters. To clarify the underlying optimization, we next emphasize the derivation of the common term $\frac{dL}{d\mathcal{P}}$ for a standard differential equation. The adjoint method requires defining a scalar functional $g(u,p)$, where $u(t,p)$ is the numerical solution to the differential equation $\frac{d\mathcal{S}(t,\mathcal{P})}{dt}=f(t,\mathcal{S},\mathcal{P})$ with $t\in[0,T]$ and initial condition $\mathcal{S}(t_{0},\mathcal{P})=\mathcal{S}_{0}$. In our experiments, we assume a known initial condition.

### v2

@@ II Thermal Runaway in Batteries
Measurements: Considering real-life constraints, only surface temperatures can be measured, and only at limited locations. For training, we take just three measurement locations, with 15 temporal measurements each, during the initial 100 minutes before the irreversible temperature rise at around 150 minutes in TR.
Analysis: Despite this limitation, Neptune accurately estimates the parameter fields ($C_{p},k_{x}$, and $k_{z}$). For the prismatic battery, the estimated parameters (Fig. 3b) show MAE of 1.88, 0.052, and 0.001, compared to reference values of 1400, 13, and 0.8, respectively. For the cylindrical battery, the scalar parameter $A_{e}$ is jointly estimated, with an estimated value of $5.19\times 10^{25}$ showing a low percentage error of 0.97% compared to the reference $A_{e}=5.4\times 10^{25}$. Similar accuracy is achieved for other parameters, with results detailed in Supplementary Note 2. Furthermore, Neptune also demonstrates robustness under noisy measurements; for the prismatic battery, results for noise levels of 1% and 10% are detailed in Table 1.
@@
Figure 4: Parameter field estimation for flow in porous media. a. A demonstration of flow phenomena, and random measurements are taken as the ground truth for model training. b. The estimated hydraulic conductivity $K$ compared to the reference $K$. c. The estimated parameter field $K$ and its estimation error compared to the reference, for different numbers of training samples. The magnitudes are normalized. d. The particle concentration predictions (state responses) at different times. The predictions align well with the ground truth, even for extrapolation at $t=20$ minutes. e. The parameter estimation performance compared to different numbers of training samples. Error bands are derived from standard deviations from multiple estimations and represent the variability in the performance. f. The response prediction performance compared to different numbers of training samples.
@@ III Flow in Porous Media
Problem Setup: A 2D computational domain of $L_{x}=1m$ and $L_{y}=0.5m$ is defined [92], as shown in Fig.4a. Using just 32 random spatial-temporal concentration measurements $u$ over 0–16 minutes, the parameter field $K$ can be robustly estimated. Since the right-hand side of equation 7 is 0, multiple solutions exist, so the estimated $K$ is normalized and compared to the normalized reference (see Fig. 4b).
@@
To demonstrate the robustness of Neptune, parameter estimation is repeated with varying sample sizes. Figure 4e shows the parameter estimation performance, while Fig. 4f displays the corresponding response prediction accuracy, with errors significantly reduced at 64 samples. Ablation studies on neural network sizes, detailed in Supplementary Note 4, further validate the model’s robustness and generalizability. For comparison with PINN-based approaches, we implemented state-of-the-art training techniques following recent best practices: residual-based adaptive sampling [95, 96] to dynamically refine collocation points in high-residual regions, 64-bit floating-point precision [97] to improve numerical stability, balanced residual decay rate (BRDR) loss weighting [98] to ensure balanced convergence across loss components, and a multi-phase training strategy with second-order optimization [24, 99] using L-BFGS-B for fine-tuning. The PINN implementation details and comparison results are provided in Supplementary Note 2.3. Additionally, a spline-based parameter fitting method was tested as a non-neural network strategy but showed lower performance than Neptune (around 2 times larger mean errors in both the parameter estimation and response prediction). The spline-based parameter fitting setups and results can be found in Supplementary Note 6.
@@
Figure 5: Parameter estimation in cell migration and proliferation. a. Parameter estimation results and comparisons. Neptune assumes all three parameters are density-dependent. The results show an increasing $\gamma$ value as density increases, while the other two parameter estimations exhibit minimal variation and can be considered as constants. This aligns well with assumptions from previous research [100, 101]. b. Density response predictions. Using the estimated parameter fields, the density response at four different timesteps is displayed. Compared to results using the reference method (PINN-SR) and the scalar parameters setting, Neptune more accurately matches the experimental data at 36-hour and 48-hour, particularly in the regions with higher density values on both sides.
@@ IV Cell Migration and Proliferation
Analysis: As shown in the first graph of Fig. 5a, Neptune estimates an increasing $\gamma$ with cell density, consistent with previous research [102, 103]. Field estimations also suggest $\lambda_{1}$ and $\lambda_{2}$ are likely scalars. We also assume space-dependent parameters for estimation: results showing parameter distributions with high (peak) values at two boundaries, demonstrating that the parameter fields are density (or state) dependent. In the third graph, while the field estimation for $\lambda_{2}$ shows large values at low densities and stabilizes as density increases, this behavior is primarily due to highly noisy experimental data in regions with low cell numbers, rather than being attributed to the nature of the parameter field itself. In contrast, PINN-SR, limited by its methodology, cannot estimate parameter fields, and its scalar estimations show slightly different combinations but remain within a reasonable range. This limitation is evident at $12h$ (first graph) and $48h$ (last graph), where it fails to capture enhanced cell spreading.
Accurate parameter field estimation excels in predicting cell density distributions over time, as shown in Fig. 5b. For such sparse and noisy data, PINN-SR achieves a median error of $6.53\times 10^{-5}$. In contrast, Neptune, assuming density-dependent parameters, achieves the highest accuracy with an error of $4.71\times 10^{-5}$. Across experimental data with varying initial cell densities [13], Neptune achieves 28-47% lower median errors. Additional results are detailed in Supplementary Note 2.
@@ V Discussion
In this paper, we presented Neptune for accurate parameter field estimation under scarce observation data, demonstrated by various PDE problems in emerging fields where physical parameters are unknown and important. Specifically, we estimated the thermal conductivity and heat capacity in coupled thermal problems, enabling accurate TR prediction. In porous media flow, spatially varying hydraulic conductivity was recovered, improving the modeling of phenomena such as contaminant transport. For cardiac EP, tissue electrical conductivity was inferred solely from voltage measurements, offering potential for detecting fibrosis and other localized pathologies. Finally, the cell migration and proliferation case demonstrates effective parameter estimation under real-world noisy and sparse data conditions. It also highlights the method’s capability in discovering governing equations and its generalizability in estimating scalar variables.
@@
In terms of utility, Neptune demonstrates its generalizability and robustness to estimate multiple parameters simultaneously across various dependencies (spatial, state, and even time), each exhibits distinct patterns. The proposed two-stage learning strategy further enhances effectiveness by handling parameters across different scales. By first estimating scalar magnitudes, it significantly reduces the search space, allowing neural networks to focus on learning local variations within the parameter field. Furthermore, we extend this discussion to demonstrate the superior performance of our strategy compared to other training approaches, including: (1) estimating only scalars, (2) directly estimating each parameter field in one step using separate neural networks; (3) estimating all parameter fields within a single neural network; and (4) jointly estimating scalar magnitudes and field variations in a single step. Detailed comparisons are provided in Supplementary Note 5. Compared to other fitting methods, Neptune significantly outperforms spline-based approaches in capturing complex parameter distributions. In addition, the use of FDM makes the method easily adaptable to a wide range of PDEs. While less effective for handling shock waves or large discontinuities, it offers favorable computational efficiency and supports gradient-based optimization.
@@
Compared to sparse identification methods like PINN-SR, Neptune models full parameter field distributions rather than scalars, achieving 27–47% improvement in forward response accuracy. This methodological shift enables more detailed modeling of the underlying physical processes, leading to improved predictive accuracy. Additionally, compared to PINN, our framework offers stronger temporal extrapolation, better adherence to physical laws, and greater adaptability to boundary and initial condition changes. It also requires less data, making it ideal when measurements are limited or costly. However, inverse problems like parameter estimation remain ill-posed and prone to local minima, especially with complex fields or limited data. These challenges can be mitigated by using informed initial guesses and incorporating prior knowledge, such as parameter magnitudes, smoothness penalties, and strategic data selection. Future work may extend the method to more general finite element frameworks and explore its application to modeling and parameter estimation in chaotic systems.
@@ VI Methods
Parameter field modeling with neural networks Neptune models the unknown parameter field $p$ as:
@@
where $\tilde{p}$ denotes a learnable scalar and $\mathcal{N}$ represents the neural network with $m$ layers and learnable weights as $\bm{\theta}$. A detailed architecture with skip connections can be referred to Fig. 6. $f$ denotes a neural layer and $\circ$ represents function composition. The input to $\mathcal{N}$ can be discretized values of space $\mathcal{X}$ or state $\mathcal{S}$, and are properly scaled (i.e., -1 to 1 from $\mathcal{X}$ or observed $\mathcal{S}$). Leaky ReLU activation functions are used in all layers except the final one, which omits activation functions. For multiple parameter fields $\tilde{\mathcal{P}}=[p_{1},p_{2},\dots]$, Neptune simply employs different scalars $\tilde{p}_{i}$ and networks $\mathcal{N}_{{\bm{\theta}}_{i}}$ for parallel modeling, where $i=1,2,\dots$
Figure 6: The schematic diagram of the two-stage parameter field estimation strategy and the neural network architecture. a. The feed-forward neural network. Left: overall architecture. Right: residual block structure. For flow and cardiac electrophysiology problems, the input to the network is the spatial coordinate $\mathcal{S}$. For battery thermal runaway and cell migration, the input is the state variable $\mathcal{X}$ (e.g., $T$ or $\rho$), with fewer layers used. The network output $\mathcal{N}_{{\bm{\theta}}}$ is scaled to the physical parameter as $\tilde{\mathcal{P}}\big(1+\mathcal{N}_{{\bm{\theta}}}\big)$. The reference value $\tilde{\mathcal{P}}$ is application-specific and determined via a two-step training strategy. b, Parameter estimation diagram. The left part illustrates the two-stage estimation for scalar and parameter fields, while the right part shows the numerical solving and optimization procedure.
@@
During the optimization, scalars $\tilde{\mathcal{P}}$ act as a physics-based prior and implicit regularizer, ensuring early physical stability of training and providing a well-scaled baseline for convenient expansion to parameter fields in the second stage. Even if $\tilde{\mathcal{P}}$ converges with high error, its combination with the PDE remains physically meaningful and stable during optimization.
@@
Compared to directly learning $\tilde{\mathcal{P}}$ with a neural network, the first-stage scalar learning significantly accelerates training, as numerically solving the PDE with parameter field modeling is several times more expensive, depending on model complexity. It also reduces the search space and greatly stabilizes the second-stage learning of $\mathcal{N}_{\bm{\theta}}$.
@@
Neural network architecture and problem-specific design The coordinate neural networks (feed-forward) in the second stage are adaptively designed, with the number of residual blocks and neurons per layer customized for each application. Specifically, in the thermal battery problem, the scalar references for thermal conductivity in different directions ($\tilde{k}_{x}$, $\tilde{k}_{z}$) and heat capacity ($\tilde{C_{p}}$) are first determined. In the cell migration problem, three scalar parameters, $\tilde{\gamma}$, $\tilde{\lambda_{1}}$, and $\tilde{\lambda_{2}}$, are similarly estimated. In both cases, the neural network models state-dependent parameter fields based on a single state variable, with a network input size of 1 (e.g., temperature $T$ or density $\rho$, rescaled approximately from -1 to 1). A compact network with 20 neurons per hidden layer and two residual blocks is used to ensure efficient training.
@@
Additionally, Supplementary Note 4 presents an ablation study evaluating the impact of neural network size on parameter estimation performance, while Supplementary Note 5 provides a detailed comparison of this training strategy against alternatives, including direct parameter field learning and using a single neural network for all parameters.
@@
During training, observations are split into mini-batches of size 16 (reduced to 8 for the cell migration problem due to limited samples). In the first stage, the loss is minimized to directly learn $\tilde{\mathcal{P}}$, with the learning rate adaptively decreasing from 1 to $10^{-2}$ until convergence, where $\tilde{\mathcal{P}}$ stabilizes without further change. Dashed lines in the figure illustrate the gradient flow during optimization. In the second stage, the loss is minimized to optimize the neural network weights $\bm{\theta}$, with a smaller learning rate ranging from $10^{-4}$ to $10^{-2}$ depending on network size. The best-performing $\bm{\theta}$, corresponding to the lowest loss value, is saved as the final trained model. The parameter field is then predicted using equation 14 with the corresponding spatial or state inputs.
@@
Importantly, Neptune differs fundamentally from traditional neural ODE approaches: instead of modeling unknown or partially known governing expressions with neural networks, it tackles complex PDEs with fully known equations, while neural networks serve only as correction terms for modeling unknown parameter fields. To clarify the underlying optimization, we next emphasize the derivation of the common term $\frac{dL}{d\mathcal{P}}$ for a standard differential equation. The adjoint method requires defining a scalar functional $g(u,p)$, where $u(t,p)$ is the numerical solution to the differential equation $\frac{d\mathcal{S}(t,\mathcal{P})}{dt}=f(t,\mathcal{S},\mathcal{P})$ with $t\in[0,T]$ and initial condition $\mathcal{S}(t_{0},\mathcal{P})=\mathcal{S}_{0}$. In our experiments, we assume a known initial condition.
@@
How Neptune outperforms PINN in inverse parameter estimation tasks? Here we demonstrate why PINN-based inverse parameter estimation fails and how our adjoint-based method fundamentally differs in its approach. This analysis uses flow in porous media as a representative example, though the moving target problem arising from simultaneous updates of parameters and state variables is general and applicable to other inverse problems where parameters $\mathcal{P}$ and state variables $\mathcal{S}$ are coupled through PDE constraints.
In the PINN formulation for inverse problems, the total loss $L_{\text{inverse}}$ consists of several components: $L_{\text{PDE}}$ represents the PDE residual losses enforcing the governing equations; $L_{\text{BC}_{i}}$ represent boundary condition losses (8 terms in total) enforcing Dirichlet and Neumann conditions on the domain boundaries. These boundary conditions are defined as follows: Hydraulic head ($h$): (1) $y=0$: Neumann flux $q=1$ (inflow); (2) $y=1$: Dirichlet $h=0$ (fixed head); (3) $x=0$: Neumann no-flux; (4) $x=0.5$: Neumann no-flux. Concentration ($u$): (5) $y=0$: Dirichlet Gaussian injection $u=\exp\left(-1600(x-0.25)^{2}\right)$; (6) $y=1$: Neumann no-flux; (7) $x=0$: Neumann no-flux; (8) $x=0.5$: Neumann no-flux. $L_{\text{IC}}$ is the initial condition loss enforcing the state at $t=0$; and $L_{\text{data}}$ is the data mismatch term comparing predicted values to observational data. The total loss is a weighted sum:
$L_{\text{inverse}}=w_{1}L_{\text{PDE}}+\sum_{i=1}^{8}w_{i+1}L_{\text{BC}_{i}}+w_{10}L_{\text{IC}}+w_{11}L_{\text{data}}$ (24)
where $w_{i}$ are loss weights.
The inverse estimation of the parameter field $\mathcal{P}$ faces a fundamental optimization challenge arising from the coupled nature of the PDE residuals. The parameter $\mathcal{P}$ (represented by neural network weights $\bm{\theta}$) is updated exclusively through the gradients of the PDE residuals, which are the only non-zero terms in $\frac{\partial L_{\text{inverse}}}{\partial\bm{\theta}}$. However, these same loss terms simultaneously update the state variables $\mathcal{S}$ through their respective gradients $\frac{\partial L_{\text{inverse}}}{\partial\mathcal{S}}$. The coupling is evident from the residual dependency $R=R(\mathcal{S},\mathcal{P})$. This creates a moving target problem during optimization: at iteration $t$, the parameter $\mathcal{P}^{(t)}$ attempts to minimize the PDE residuals $R(\mathcal{S}^{(t)},\mathcal{P}^{(t)})$ based on the current state $\mathcal{S}^{(t)}$, but these state variables are concurrently being updated to $\mathcal{S}^{(t+1)}$ by the same residual terms. As a result, the optimization landscape for $\mathcal{P}$ shifts with each update, preventing stable convergence to the true parameter values. In contrast to the PINN approach, Neptune circumvents this optimization conflict by decoupling parameter estimation from state variable learning. The forward problem is solved using a numerical solver, which enforces a deterministic relationship $\mathcal{S}=\text{Solver}(\mathcal{P})$ between the parameter and the state variable. Critically, during optimization, only the neural network weights $\bm{\theta}$ are updated; the state variable $\mathcal{S}$ is not a free optimization variable but is uniquely determined by $\mathcal{P}$ through the differentiable PDE solver at each iteration. This direct, one-way coupling eliminates the moving target problem: at iteration $t$, the parameter $\mathcal{P}^{(t)}$ minimizes the loss based on $\mathcal{S}^{(t)}$, and the subsequent state $\mathcal{S}^{(t+1)}$ is fully determined by the updated parameter, ensuring a stable optimization landscape that enables reliable convergence to the true parameter values.
@@
Data availability All synthetic training data are generated on-the-fly by our provided scripts. The experimental cell migration data, originally published in [104], are also packaged within our code repository. Additional information can be found in the Code availability section.
Code availability The source code used to generate the results presented in this study is provided as supplementary file for the peer-review process. The main algorithm is implemented in Julia, while baseline methods including PINNs and neural operators are implemented in Python (PyTorch). Upon acceptance, the code will be hosted on a public GitHub repository.